OpenStudy (anonymous):

Which is a constant in the expression 6j +1 + 2k?

OpenStudy (anonymous):

help plz

OpenStudy (ericwise):

1, the other numbers multiplied by a variable are coefficients

OpenStudy (anonymous):

A variable is something for which the value can vary. In this example both j and k are variables. This is because if we were to write, for example: j = 1 + 1 = 2 then j has a different value than if we wrote j = 1 + 2 = 3. The value of j can change from one question (or instance) to the next. "6", "1" and "2" above are all constants however. Their value never changes.

OpenStudy (ericwise):

The constant is "1"

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

boiii